OpenStudy (anonymous):

Solve the logarithmic equation. Express your solutions in exact form only. Please show all of your work. log3(x+10)+log3(x-3)-2=log3(x-2)

OpenStudy (anonymous):

\[ \log_{3}{(x+10)}+log_{3}{(x-3)}-2= log_{3}{(x-2)}\] \[ 3^{\log_{3}{(x+10)}+log_{3}{(x-3)}-2}= 3^{log_{3}{(x-2)}}\] \[ ((x+10)*(x-3))*3^{-2}= (x-2) \] \[ x^{2}+7x-30= 3^{2}(x-2) \] \[ x^{2}-2x-48=0 \] \[(x-8)(x+6)=0\] \[x=\{8, -6\}\] Plug both answers into the original equation to check for extraneous roots. I skipped a few steps, but am sure you can fill them in, remember the rules for logs when going through this.

OpenStudy (anonymous):

Thank you. :)